Fermi National Laboratory
 

Accelerator Update for October 23, 2006 - October 25, 2006

 

Monday October 23

The day shift began with the Tevatron (TeV) turn on continuing, with the Antiproton Source (Pbar) stacking (60.1E10), with Operations monitoring the Recycler stash (109.9E10), with Linac, Booster, and the Main Injector (MI) operating normally, with MiniBooNE and NuMI taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for colliding beam.

TeV experts began start-up studies at 8:31 AM.

Operations injected 14 bunches of protons into the TeV and then put them through a wet squeeze at 10:12 AM.

At 10:35 AM, the TeV beam aborted due to a quench at sector B4. The cryo system at sector B4 went bad and this caused the abort.

At 11:55 AM, cryo system technicians reported that the inverter box for the TeV sector B4 cold compressor failed during the quench. They began work to replace the box.

Operations prepared D0 and CDF for access at 12:06 PM.

Operations put the TeV through a dry squeeze at 3:44 PM.

Operations put the TeV into shot setup at 5:01 PM.

Operations established store 5019 at 7:12 PM with an initial luminosity of 102.03E30.

The Recycler completed a stash of antiprotons around 7:46 PM.

At 7:55 PM, Operations reported loosing beam permits for the P1 and P2 transfer lines. An expert suggested a timing change and this fixed the problem. The NuMI beam permit also went down, but due to a horn ground fault. Operators reset the horn per expert instruction.

Tuesday October 24

The midnight shift began with Operations monitoring store 5019 (55.42E30) and the stash 129.4E10), with Pbar stacking (20.3E10), and with NuMI and MiniBooNE taking beam.

The Recycler completed a stash around 2:31 AM.

At 2:46 AM, the TeV suffered a quench at B4. The B4 cryo system went unstable again.

At 3:39 AM, the Cryo Coordinator reported that the B4 cold compressor needed to be replaced.

The Recycler completed a stash at 6:03 AM.

Cryo techs replaced the B4 cold compressor at 6:37 AM. System cooldown began.

The Recycler completed a stash around 9:35 AM.

Cryo system techs found and replaced a bad cable on the on the B4 system at 10:33 AM. This was the cause of the B4 problems.

Operations put the TeV through a dry squeeze at 11:58 AM.

Operations put the TeV into shot setup at 12:28 PM.

Operations established store 5021 at 3 PM with an initial luminosity of 175.6E30.

Recycler experts began a study around 3:04 PM.

Around 4:39 PM, store 5021 aborted due to a safety system glitch. The TeV quenched at E1 and D1. The F sector audio permit dropped and took down several critical device controllers. This failure held off beam to NuMI, MiniBooNE, and Pbar.

Interlock techs reported at 5:53 PM that they had replaced a F sector safety system board.

Operations reported at 5:56 PM that the B4 cryo system had recovered.

Operations resumed sending beam to NuMI, MiniBooNE, and Pbar at 6:13 PM.

Operations injected one bunch of protons into the TeV at 9:34 PM. Experts began a study.

The TeV study finished at 11:24 PM.

The Recycler completed a stash at 11:36 PM.

Operations put the TeV into a dry squeeze at 11:53 PM.

Wednesday October 25

The midnight shift began with the TeV preparing for a shot setup, with Operations monitoring the stash (154.2E10), with Pbar stacking (23.9E10), and with NuMI and MiniBooNE taking beam.

Operations put the TeV into shot setup at 12:19 AM.

Operations established store 5026 at 2:23 AM with an initial luminosity of 128.9E30.

At 4:03 AM, Operations reported a drop in luminosity, but no beam was lost. Later, a TeV expert attributed the drop to a D17 separator spark.

The Recycler completed a stash at 5:41 AM.

The Plans for Wednesday

The plan for today is stack and stash, and allow some studies. The Run Coordinator has not decided when the next store would go in, but possibly late on the evening shift or early on the midnight shift.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/25/2006    email Fermilab